1. (4 points) Suppose that the markup of goods over unit costs is 20% and that that the wage-setting equation is W = P(1 - u)Please help with C and D. Thank You.
Answer #1

A. wage setting equation is w= P(1-u)

Therefore price setting equation would be P= (1+m)w [where m is the increase price over cost]

m=20% which is .2

thus,

P=(1+0.2)W

W/P = 1/1.2

= 0.8333

B. From the wage settings, we can calculate the unemployment rate

Un= 1-W/P

= 1-0.833

= 0.1667

Thus unemployment rate is = 16.67%

C. If the mark up prices over costs increases to 25%,

W/P = 1/1.25 [refer point A to get derivation]

= 0.8

Therefore unemployment rate is calculated by 1-W/P

= 1-0.8

=0.2

thus the unemployment rate rises to 20%. This is because the increase in the markup lowers the real wage. And from the wage-setting equation, the unemployment rate must rise for the real wage to fall. Also an increase in the markup implies more market power for firms, and therefore less production, since firms will use their market power to increase the price of goods by reducing supply.

D. Firms are able to change their mark ups because of the presence of competition and that it is not a perfectly competitive market. They can adjust their costs, supply and all variable factors to bring about changes in the price over cost to earn more. In this type of a situation the prices are never fixed.

On the other hand, if it was a perfectly competitive market, the prices would remain fixed and constant for all the players in the market.